# secReferences in the Online Appendix

### v5

@@ Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ strategies and how they evolve over time. Data from 2,500 campaigns show that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer backlash and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
@@
Keywords: Non-Governmental Organizations, campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential actors shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017). However, their actions are difficult to observe, and evidence on their strategic incentives remains limited (Yaziji and Doh, 2009). As a result, NGOs are often viewed either as principled actors promoting social good (Daubanes and Rochet, 2019) or as dysfunctional entities vulnerable to corporate capture (Bertrand et al., 2020).
This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value. We focus on campaign timing—one of the few observable aspects of NGO activity, which is far more opaque than that of for-profit firms. [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base. Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored, especially outside developing-countries contexts (e.g., Deserranno et al., 2025). While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior. We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize exposure by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals. However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly. Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008). However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns. In contrast, past non-AGM campaigns have no influence on proposal success. NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time. This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant. Our main endogenous variable, therefore, is NGO campaign timing, which we explain by reference to each NGO’s prior campaign experience: more experienced NGOs campaign earlier, trading visibility for impact. While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience or campaign timing. This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues. NGOs also engage firms through partnerships and other channels, but since such activities are voluntarily disclosed—and thus subject to selection biases (Burger and Owens, 2010)—we rather focus on campaign timing, which must be public to be effective and, thus, observable to researchers.
Our findings open new avenues for research, as future work may empirically explore how NGOs choose between campaigning and other activist tactics or whether these different approaches act as complements or substitutes. The exogenous timing of AGMs provides a useful identification strategy for such analyses, as it holds the target firm’s strategic choices constant, allowing researchers to isolate NGOs’ choices and their incentives.
Our results reveal a novel tradeoff between exposure and direct impact as a key driver of the societal consequences of NGO activism (Briscoe and Gupta, 2016). Unlike our framework, much of the existing literature assumes that activists primarily pursue social objectives, either stemming from prosocial behavior (Besley and Ghatak, 2005; Bénabou and Tirole, 2006) or moral preferences (Baron, 2010), without discussing their incentives.
At the same time, several studies highlight instances in which NGOs themselves are subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). Firms may also strategically leverage their relationships with NGOs—through donations or partnerships—to advance political agendas (Bertrand et al., 2020), or to restore consumer and investor trust following reputational damage (Akey et al., 2021).
Our main contribution in this literature is to show that NGOs are not passive actors. Rather, their tactics evolve over time, shaping firms and their stakeholders’ responses in different ways depending on the strategies they employ (e.g., Hendricks and Sorensen, 2009; Cabral and Hortacsu, 2010). Thus, understanding NGOs’ incentives is essential to learn why firms internalize their externalities.
We also contribute to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear (e.g., Hawn and Ioannou, 2016). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010). However, other agents may play a pivotal role. Among them, we highlight that NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Koenig and Poncet, 2019; Fioretti et al., 2023; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and concludes.
2 Data
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. Those single-firm campaigns represent 93% of the original sample. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms by market capitalization. For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. [^3] Evidence shows that shareholders generally agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
Lastly, we source firm-level news data from Ravenpack. RavenPack is a data analytics platform that transforms unstructured news content into structured data, offering sentiment analysis and event detection. The data allows us to count the news related to every firm headquartered in the US. [^4] We only retain news with a Relevance Score of at least 75%, which ensures the relevance of the news for the firm. We include only US-listed firms from 2010 to 2020. The sample includes 5,212 US firms.
We link these three datasets using standard identifiers, including year, firms’ CUSIP, and ISIN numbers. Appendix Table A1 presents summary statistics for NGO-related variables (Panel a). On average, NGOs launch 3.40 campaigns during the sample period, but the distribution is skewed, with over 50% of NGOs initiating only one campaign. Most campaigns focus on social issues rather than environmental ones. Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
The following three sections examine an NGO’s payoffs in connection with the timing of its campaigns, and the reactions of central stakeholders. We present the key takeaways from each analysis at the end of each section. Before that, we detail our empirical approach.
3 AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B2 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^5] Results remain robust excluding the covid period (2020–2021). While external events may influence NGOs’ campaign decisions, they do not affect AGM timing. Thus, AGMs allow us to isolate NGOs’ strategic trade-off from firms’ responses.
Target selection. While our analysis centers on campaign timing—leaving NGOs’ target selection to future research—Appendix Table A2 shows that NGOs campaigning on AGM dates do not select firms that differ in observable characteristics (market capitalization, assets, sales, media coverage, or environmental and social scores) from those targeted at other times, and campaign sentiment is likewise comparable across both groups. Moreover, Table A1 demonstrates that target choice is independent of prior campaign experience: Panel (c) reveals no significant differences in market cap, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs, and Panel (d) finds no correlation between firm characteristics and NGOs’ past campaigns. However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
4 Timing Campaigns for Media Exposure
NGOs are substantially more likely to launch their campaigns around AGM dates, on average. According to Figure 1(a), NGO campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100.666Appendix Figure B3 confirms this finding showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$) (Cattaneo et al., 2020).
@@ 4 Timing Campaigns for Media Exposure
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack. Calendar weeks are re-centered around the firm’s AGM.
This finding could be related to the heightened media coverage during AGMs (Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
@@
$\text{News}_{n,w,y}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}{\{M_{n,w,y}=\tau\}}\,\cdot\,\text{NGO Campaign}_{n,w,y}\right)$ (1)
$+\mathbf{X}_{n,w,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y}\,\,,$
where “$\text{News}_{n,w,y}$” is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The independent variables include a set of dummies “$\mathbf{1}\{M_{n,w,y}=\tau\}$” indicating the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.777 The analysis aligns the calendar year with the corporate year, starting five months before the AGM month (month 0) and ending five months later, with months -5 and +5 slightly longer. Since SEC regulations make AGM months stable over time (Section 3), this relabeling does not alter the distribution of months by firms or double-count adjacent years. We interact these event-study indicators, with the variable “$\text{NGO Campaign}_{n,m,y}$,” an indicator variable equal to 1 if firm $n$ experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. We set the reference month to four months before the AGM, following SEC rules that require the AGM date to be set at least four months in advance.
Figure 2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Standard errors are clustered at the firm level.
The vector “$\mathbf{X}_{n,w,y}$” includes “$\mathbf{1}{\{M_{n,w,y}=\tau\}}$” and “$\text{NGO Campaign}_{n,m,y}$” separately, and controls for (i) the occurrence of negative news and (ii) all E&S news over the previous four weeks to account for a firm’s recent media coverage. [^8] Adding controls for the current count of E&S news to account for NGOs targeting AGMs of high-coverage firms does not qualitatively change the results. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-week-by-year fixed effects ($\rho_{w,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients. Negative news follow a bell-shaped pattern around the AGM, peaking significantly above zero in the AGM month, highlighting the benefits of targeting AGMs. [^9] In section 6.2, we show that the spike on AGM date is particularly strong for an NGO’s first few campaigns but it fades away for later ones (Appendix Figure B4). To quantify the magnitude of the AGM peak, an NGO campaign launched in the AGM week doubles the probability of a negative E&S news compared to a campaign initiated four months earlier, on average. After the AGM, coverage returns to baseline levels, indicating that NGOs attract more media attention when campaigning closer to the AGM.
@@
Media coverage of NGO campaigns is substantially greater during the target’s AGM.
Importantly, we rule out the hypothesis that NGOs select more negative causes when targeting AGMs. Without an adequate ranking of causes by negativity, we estimate (1) with firm $n$’s average E&S news sentiment in week $w$ as the dependent variable. [^10] RavenPack scores the tone of a firm’s news coverage between -1 (adverse) and 1 (favorable). Examples of positive and negative news can be found in Appendix Figure B5. The estimated coefficients show that campaigns launched during AGM months are no different than other campaigns (Appendix Figure B6).

### v6

# NGO Activism:

@@ Abstract
We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs). Using data on 2,500 U.S. campaigns, we find that NGOs frequently launch campaigns on the targeted firm’s AGM date, which boosts media exposure and triggers future shareholder proposals but does not sway current AGM votes or affect sales. As NGOs gain experience and reputational capital, they shift to earlier pre-AGM campaigns that increase shareholder support at the AGM and lead to improvements in corporate behavior. These patterns reveal a strategic trade-off between maximizing exposure and achieving direct influence.
@@
Keywords: Non-Governmental Organizations, NGO campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
@@ 1 Introduction
Non-governmental organizations (NGOs) are increasingly prominent actors in shaping firms’ environmental and social policies (Baron, 2001; Egorov and Harstad, 2017; Yaziji and Doh, 2009). Yet, their strategic choices remain poorly understood. Much of the literature treats NGOs as either principled advocates motivated by social good (Besley and Ghatak, 2005; Bénabou and Tirole, 2006; Daubanes and Rochet, 2019) or as dysfunctional entities prone to corporate capture and strategic manipulation (Grossman and Helpman, 1994; Bertrand et al., 2020; Akey et al., 2021). We argue and demonstrate empirically that NGOs adapt their strategies in response to incentives, shaping which stakeholders they engage and ultimately affecting the social and environmental performance of targeted firms.
To study how NGOs respond to incentives and their implications for firms, we focus on a setting where firms are especially exposed to outside pressure: the Annual General Meeting (AGM). AGMs are predictable focal points at which executives present results, shareholders vote on proposals, and media attention peaks (Fioretti et al., 2025). Figure 1 shows that NGO campaigns cluster heavily around AGM weeks: incidence is four to six times higher than in non-AGM weeks (Panel a), and media coverage spikes simultaneously (Panel b). These striking patterns raise a central question: do NGOs target AGMs mainly to maximize exposure, or do they also use campaigns to exert influence over corporate outcomes?
@@
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,544 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack.
Examining campaign timing reveals that less experienced NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020). These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents. For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025). Indeed, we find little evidence that other stakeholders are moved by these campaigns. For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves. Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively inexperienced. These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times. This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
As NGOs accumulate visibility through AGM-date campaigns, their strategic calculus changes. Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes. By contrast, campaigns launched in the months leading up to the AGM substantially raise the likelihood that related shareholder proposals succeed. Intervening one to three months before the meeting boosts average shareholder support by 8%, while even earlier campaigns often prompt firms to settle before proposals reach a vote, leading to measurable improvements in environmental and social performance.
Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns forward in time as they gain experience and reputational capital. First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions. This pattern is especially pronounced among NGOs with stronger reputations, proxied by higher search visibility, which campaign several months ahead of the AGM rather than on the day itself.
Repeated targeting of AGMs also plays a distinct role. By returning to these focal moments, NGOs build credibility and reputational standing, which in turn increases the effectiveness of later campaigns. Yet such returns are not indefinite: once credibility is established, additional AGM-day campaigns yield diminishing benefits. Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—shows how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for firms’ social and environmental impact.
In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As NGOs accumulate reputational capital (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their strategy shifts: campaigns are launched months before the AGM, when they can still sway shareholder support and discipline managers.
This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022). More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
Prior studies highlight the role of shareholders influencing firms via voice through voting or exit via divestments (e.g., Cuñat et al., 2016; Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), and of consumers through willingness to pay (Gneezy et al., 2010). We instead show that NGOs act as coordination mechanisms, enabling diverse stakeholders to collectively discipline firms (Koenig and Poncet, 2019; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023). In doing so, NGOs expand the set of channels through which civil society shapes corporate conduct.
Finally, our methodology relates to the literature on the strategic timing of actions to exploit fluctuations in media attention. Prior work shows that politicians and governments time executive orders, military operations, and even disaster relief to coincide with or avoid competing news (Eisensee and Strömberg, 2007; Durante and Zhuravskaya, 2018; Djourelova and Durante, 2022), and that social movements strategically leverage focal dates to amplify visibility (Madestam et al., 2013). We show that NGOs adopt similar tactics in the corporate domain.
The paper is organized as follows. Section 2 outlines the data, and Section 3 provides background on AGM dates. Sections 4 and 5 examine NGO campaigns as tools to gain media exposure or to influence firms through key stakeholders, respectively. Section 6 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals. Section 7 discusses the results and concludes.
2 Data
Our dataset integrates four sources: NGO campaigns, online attention, shareholder proposals, and firm-level news.
First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data record the campaign’s public release date, cause (e.g., “GMOs in food”), the targeted company and its prominence, and sentiment (ranging from –2 to +2, typically negative). We focus on single-firm, negative-sentiment campaigns where the firm appears in the headline. [^1] Since our analysis links campaign timing with a firm’s AGM date, we exclude multi-firm campaigns as their AGM dates may differ. Single-firm campaigns represent 93% of the original sample. The final dataset spans 2,512 campaigns launched by 731 NGOs against 523 firms.
Second, we collect monthly Google Trends scores for all NGOs in our sample from 2010 to 2022. Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches (within-NGO) or (ii) relative to a benchmark NGO in the same month (between-NGOs). We use Greenpeace—the most searched NGO—as the benchmark for comparability across organizations.
Third, we use data from Institutional Shareholder Services (ISS) on proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms. We select 4,483 environmental and social (E&S) proposals filed between 2010 and 2021 (837 firms in total). About one-third of these proposals are successful, either withdrawn before the AGM (after firm commitments) or passed in a vote. Both the number of E&S proposals and their success rate increased steadily over the period.
Lastly, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging. We count news items related to US-listed firms from 2010 to 2020, retaining only items with a Relevance Score of at least 75%. The final sample includes 5,212 firms.
We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year). Appendix Table A1 reports descriptive statistics. Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once. Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign. Panel (b) documents substantial variation across targeted firms in size and news coverage. Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations. 53% of the campaigns are on social causes, with the remaining being environmental. We return to the remaining panels of Table A1 in Section 4, after first providing background on AGM dates.
3 Institutional Background on AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B1 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^2] Results remain robust excluding the covid period (2020–2021).
While external events may shape NGOs’ incentives to campaign, they do not alter the timing of AGMs. This stability makes AGMs a useful benchmark: NGOs can choose to react immediately to a negative event concerning a firm, or alternatively wait and align their campaign with its AGM, when the firm is most exposed to scrutiny, as shown in Panel (b) of Figure 1, which shows that news coverage spikes in AGM weeks. Consistent with this view, we show in the Appendix that AGM-timed campaigns generate more negative news coverage than those launched at other times (Appendix Figure B2). Taken together, these patterns highlight why AGMs provide a natural focal point to study the strategic trade-off between maximizing visibility and exerting direct influence.
4 Timing Campaigns for Visibility
Having shown that NGO campaigns cluster around AGM dates and that they increase a target’s negative media coverage, we next ask whether such timing enhances NGOs’ visibility. We use Google search activity as a proxy for NGO visibility and estimate event-study regressions comparing campaigns launched on the AGM date with those launched at other times:
@@ 4 Timing Campaigns for Visibility
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$
$\quad+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
Because NGOs typically campaign multiple times, we follow Baker et al. (2022a) and stack all campaign events. Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$. We then analyze variation in the Google Trends score for NGO $g$ within a five‐month window around the start of $g$’s campaign; this window may spill into years y1y-1 or $y+1$ if the campaign occurs early or late in year $y$.
The main explanatory variables are event-time dummies, $\mathbf{1}_{\{M_{c,m,y}=\tau\}}$, equal to one if month $m$ is exactly $\tau$ months from campaign $c$, with $\tau\in[-5,5]$. The omitted category is the month immediately preceding the campaign ($\tau=-1$). These indicators are interacted with $\text{AGM Campaign}_{c}$, a dummy equal to one if campaign $c$ was launched on the target firm’s AGM date. Thus $\beta^{AGM}_{\tau}$ estimates the increase in visibility for NGO campaigning on an AGM date, over the direct effect of campaigning at any date, $\beta^{Non-AGM}$.
The remaining terms include fixed effects at the NGO–firm–year level ($\alpha_{c,y}$), which absorb all variation across NGO-firm pairs over years, as well as month-by-year fixed effects ($\alpha_{m,y}$). Our main coefficients of interest ($\beta^{AGM}_{\tau}$) are therefore identified from within-pair variation across months. Standard errors are clustered at the NGO level.
Under the hypothesis that AGM campaigns affect the visibility of experienced and inexperienced NGOs differently, we estimate (1) separately for an NGO’s first few campaigns and for later ones. Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and its subsequent campaigns (red triangles). The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months. For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established. These patterns suggest that inexperienced NGOs rely more heavily on AGM timing to build name recognition, whereas established NGOs gain less from repeating the same tactic.
Figure 2: Increased NGO visibility around an AGM campaign
(a) The first AGM campaigns bring visibility
(b) No evidence of Selection
Note: Panel (a) plots the estimated $\beta^{AGM}$ coefficients from (1). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level. Panel (b) reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects. Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
We show that these results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with much larger gains for “unexperienced” NGOs.
Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure. Panel (b) of Figure 2 addresses this concern by regressing an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the cause of the campaign, with year fixed effects and firm-clustered standard errors. [^3] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), the firm’s Environmental and Social scores (Env. score, Soc. score), how negative the campaign is towards the firm (Campaign sentiment), and an indicator for the campaign cause being related to an environmental topic (Env. campaign). Note that no coefficient is significantly different from zero. We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
Further evidence on selection based on NGO experience comes from Table A1. Panel (c) compares firms targeted by NGOs with high versus low campaign experience and finds no significant differences across the same variables studied in Panel (b) of Figure 2. Panel (d) corroborates this by regressing firm characteristics on NGO experience (i.e., number of past campaigns), showing no correlation once fixed effects are included. The only systematic trend is that NGOs increasingly focus on environmental rather than social issues in later years, reflecting their rising importance.
Thus, the evidence presented so far points to:
@@
Gains in visibility are greater for unexperienced NGOs campaigning on the target’s AGM day than for experienced NGOs.
For clarity, each of the next sections highlights its main findings in a similar manner.